# Automatically Interpreting Millions of Features in Large Language Models

### v1

@@ A.7.3 Baselines
Random SAE. We experiment with a random top-k SAE with $k=50$, roughly the average sparsity of the gemma SAEs. The encoder is initialized with 131,072 spherically uniform unit-norm features, and the decoder is initialized to its transpose. We use a random SAE with TopK activations because we need sparsity for our sampling procedure to work properly.
Random explanations. For each layer and target KL value, we compute a random explanation baseline where we shuffle the explanations across features.
@@ A.7.4 Intervention interpretability results
Figure A6: Average intervention score vs layer. SAE features in later layers have more explainable effects on output. Random features, however, have uninterpretable effects on output, even at late layers.

### v2

@@ A.7.3 Baselines
Random SAE. We experiment with a random top-k SAE with $k=50$, roughly the average sparsity of the gemma SAEs. The encoder is initialized with 131,072 spherically uniform unit-norm latents, and the decoder is initialized to its transpose. We use a random SAE with TopK activations because we need sparsity for our sampling procedure to work properly.
Random interpretations. For each layer and target KL value, we compute a random interpretation baseline where we shuffle the interpretations across latents.
@@ A.7.4 Intervention interpretability results
Figure A6: Average intervention score vs layer. SAE latents in later layers have more explainable effects on output. Random latents, however, have uninterpretable effects on output, even at late layers.